OpenStudy (anonymous):

Find solutions for 2cos^2(theta)-sin(theta) =1 from [0,2pi)

OpenStudy (anonymous):

method is to rewrite \[cos^2(t) = (1-sin^2(t)\]

OpenStudy (anonymous):

\[2(1-sin^2(t))-sin(t)=1\] \[2-2sin^2(t)-sin(t)=1\] \[2sin^2(t)+sin(t)-1=0\]

OpenStudy (anonymous):

now you have a quadratic equation in \[sin(t)\]

OpenStudy (anonymous):

like solving \[2x^2+x-1=0\]

OpenStudy (anonymous):

where \[x=sin(t)\] \[2x^2+x-1=0\] \[(2x-1)(x+1)=0\] \[x=-1\] or \[x=\frac{1}{2}\]

OpenStudy (anonymous):

so now solve \[sin(t)=-1\] means \[t=\frac{3\pi}{2}\]

OpenStudy (anonymous):

\[sin(t)=\frac{1}{2}\] \[t=\frac{\pi}{6}\] or \[t=\frac{5\pi}{6}\]

OpenStudy (anonymous):

done!

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

welcome!